xx TimerInfo

TimerInfo

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11clock 10Scope: privatex01  STRING

11Description: Which clock to use

11Range10 x01 Default: all
11.*  10Anything

11

11every 10Scope: privatex01  INT

11Description: How often to print information

11Range10 x01 Default: (none)
110:* 10Positive

11

11timer 10Scope: privatex01  STRING

11Description: Which timer to use

11Range10 x01 Default: all
11.*  10Anything

11

7 Interfaces

General Implements: timerinfo

8 Schedule This section lists all the variables which are assigned storage by thorn CactusExamples/TimerInfo. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE Scheduled Functions

CCTKxPOSTSTEP x timerinfo x print information from cactus timers x

11x11Language:11c
11 x 11Type: 11function